# Q-VLM: Post-training Quantization for Large Vision-Language Models

### v1

@@ 2.2 Post-training Quantization
Network quantization substitutes full-precision tensors with low-precision values and replaces multiply-accumulate operations with integer arithmetics, which significantly reduces the storage and computational cost of neural networks. Traditional quantization-aware training (QAT) methods [5, 30] need to finetune network weights with the full training set for rounding, which is less practical since the data and resources for training may not be accessible for most users. Recently, post-training quantization (PTQ) [10, 31, 39, 9] has aroused extensive interest, which leverages a small calibration set to search for the optimal threshold in rounding functions with significantly reduced data demand and optimization cost. Choukroun et al.  [6] minimized the $l_{2}$ distance between quantized and full-precision tensors to mitigate evident task performance degradation, while Zhao et al.  [46] duplicated channels with outliers and halved their values to reduce clipping loss without amplifying rounding errors. Liu et al.  [31] preserved relative ranking orders of self-attention in vision transformers to mitigate information loss during post-training quantization and explored a mixed-precision quantization strategy based on the nuclear norm of attention maps and features. Zero-shot PTQ further extends the boundaries for efficiently quantizing neural networks without real image data. Cai et al.  [4] optimized pixel values of generated images to align sample batch statistics with those of batch normalization (BN) layers in full-precision networks. Li et al.  [25] extended the PTQ framework to transformer architectures by diversifying self-attention across different patches using patch similarity metrics. Meanwhile, deploying PTQ to large language models (LLMs) dynamically search the optimal rounding function for each input sample instead of applying a learnable one, because the activation distributes very significantly across different samples in large models. LLM.int8() [7], SmoothQuant [41] and ZeroQuant [42] handled activation outliers to achieve accurate quantization function learning by eliminating the extreme values with equivalent transformation, yet encountered challenges in effectively scaling to extremely large models due to the unbearable computational costs. Furthermore, GPTQ [11], AWQ [27], and QLoRA [8] deployed low-precision quantization on weight quantization to further reduce the computational complexity. However, these methods employ layer-wise searching strategy to search the rounding functions sequentially, which deviates from the optimal ones due to the lack of cross-block dependency.
@@ 3.2 Mining Cross-layer Dependency for LVLM Quantization
Figure 2: The correlation between discretization error difference (DED) and the activation entropy.
@@
Explicitly evaluating the cross-layer dependency requires multiple forward pass of LVLMs for given input, where the correlation for discretization errors is calculated from data statistics. To avoid extremely high cost, we have to explore an efficient proxy to approximate the cross-layer dependency. We leverage Information Entropy of the activation to judge the sensitive layers with homogeneous distribution. For sensitive layers, the noise in the former layer caused by the deviation from the global optimal value usually leads to higher deviation for the current layer, so that jointly search these layers decrease the block output quantization errors. As a result, discretization error difference (DED) between layer-wise search and joint search is obvious for sensitive layers with high entropy. In order to reduce the deviation from the optimal rounding points, we should search the quantization function by joint consideration of the current layer and the former one with strong dependency.

### v2

@@ 2.2 Post-training Quantization
Network quantization substitutes full-precision tensors with low-precision values and replaces multiply-accumulate operations with integer arithmetics, which significantly reduces the storage and computational cost of neural networks. Traditional quantization-aware training (QAT) methods [6, 32] need to finetune network weights with the full training set for rounding, which is less practical since the data and resources for training may not be accessible for most users. Recently, post-training quantization (PTQ) [11, 33, 44, 10, 42, 41] has aroused extensive interest, which leverages a small calibration set to search for the optimal threshold in rounding functions with significantly reduced data demand and optimization cost. Choukroun et al.  [7] minimized the $l_{2}$ distance between quantized and full-precision tensors to mitigate evident task performance degradation, while Zhao et al.  [52] duplicated channels with outliers and halved their values to reduce clipping loss without amplifying rounding errors. Liu et al.  [33] preserved relative ranking orders of self-attention in vision transformers to mitigate information loss during post-training quantization and explored a mixed-precision quantization strategy based on the nuclear norm of attention maps and features. Zero-shot PTQ further extends the boundaries for efficiently quantizing neural networks without real image data. Cai et al.  [5] optimized pixel values of generated images to align sample batch statistics with those of batch normalization (BN) layers in full-precision networks. Li et al.  [27] extended the PTQ framework to transformer architectures by diversifying self-attention across different patches using patch similarity metrics. Meanwhile, deploying PTQ to large language models (LLMs) dynamically search the optimal rounding function for each input sample instead of applying a learnable one, because the activation distributes very significantly across different samples in large models. LLM.int8() [8], SmoothQuant [46] and ZeroQuant [48] handled activation outliers to achieve accurate quantization function learning by eliminating the extreme values with equivalent transformation, yet encountered challenges in effectively scaling to extremely large models due to the unbearable computational costs. Furthermore, GPTQ [12], AWQ [29], and QLoRA [9] deployed low-precision quantization on weight quantization to further reduce the computational complexity. However, these methods employ layer-wise searching strategy to search the rounding functions sequentially, which deviates from the optimal ones due to the lack of cross-block dependency.
@@ 3.2 Mining Cross-layer Dependency for LVLM Quantization
Figure 2: The correlation between discretization error difference (DED) and the activation entropy in 15th layer.
@@
Explicitly evaluating the cross-layer dependency requires multiple forward pass of LVLMs for given input, where the correlation for discretization errors is calculated from data statistics. To avoid extremely high cost, we have to explore an efficient proxy to approximate the cross-layer dependency. We leverage Information Entropy of the activation to judge the sensitive layers with homogeneous distribution[43]. For sensitive layers, the noise in the former layer caused by the deviation from the global optimal value usually leads to higher deviation for the current layer, so that jointly search these layers decrease the block output quantization errors. As a result, discretization error difference (DED) between layer-wise search and joint search is obvious for sensitive layers with high entropy. In order to reduce the deviation from the optimal rounding points, we should search the quantization function by joint consideration of the current layer and the former one with strong dependency.
@@ 4.2 Ablation Study
Figure 4: Visual reasoning examples from LLaVA-13B model. Q-VLM improves over the AWQ baseline for W4A4 quantization, reducing quantization errors and providing more reasonable answers. We color the text to show the correct or wrong responses.
Visualization reasoning examples: We further provide qualitative visual reasoning example of the LLaVA-v1.3-13B model in Figure 4 from ScienceQA dataset. Q-VLM improves the responses compared to AWQ baseline for W4A4 quantization setting, leading to more reasonable and particular answers for vision-language question pairs. In this first example, Q-VLM correctly answers the question about the highlighted ocean, while AWQ appears to have limited comprehension of the image information. For the second example, AWQ under W4A4 lost substantial information to produce sound reasoning about whether Fromia monilis cells make their own food, while Q-VLM answered correctly and even produced a table for precise classification about the six broad organisms called kingdom. Q-VLM improves the visual reasoning ability of LVLMs by reducing factual errors in the responses.
@@ Acknowledgement
Appendix A Why leveraging entropy as proxy:
The benefit and motivation for using entropy rather than quantization errors as a proxy for block-wise searches lie in several key considerations. We analyze that larger entropy indicates more homogeneous data distribution, which is a well-established principle in information theory. Consequently, DED and activation entropy are strongly correlated with an value of 0.97. However, greater quantization error does not necessarily imply more homogeneous data distribution and does not show a positive correlation with DED, having an value of 0.81, which is empirically verified in the figure 5.
Meanwhile, the search cost of quantization errors doubles compared with entropy as a proxy, as the calculation of quantization errors requires multiple forward passes for both the FP model and the quantized model. The weak correlation and the unbearable search cost render quantization error unsuitable as a metric for measuring cross-layer dependency.
Furthermore, we conducted experiments comparing the proxy effectiveness of quantization error and entropy across different models under various bitwidths in Table 5. Entropy outperformed quantization errors by a significant margin (78.66 vs. 77.97), showing a strong cross-layer dependency within each block. This allowed us to achieve optimal block partitioning by mining the cross-layer dependency.
Model Dependency Proxy W6A6 W4A4
Accuracy Search Cost Accuracy Search Cost
LLaVA-7B Quantization Errors 88.59 43.7 77.97 41.2
Entropy 88.95 26.9 78.66 25.9
LLaVA-13B Quantization Errors 88.96 58.6 78.82 54.2
Entropy 89.04 33.1 79.89 31.8
Table 5: Comparisons with different proxy for mining cross-layer dependency for LLaVA-v1.3 models in ScienceQA dataset across bitwidth setting.
Appendix B Performance on more baseline methods
We have extended our experiments to an additional baseline method ZeroQuant-V2[47] and compared it against our proposed methods in Table 6. ZeroQuant-V2 leverages per-token quantization with different rounding functions to minimizing activation discretization errors. However, ignoring cross-layer dependency of discretization errors fails to acquire the optimal rounding strategy with severe outliers under low bitwidth and degrades the performance significantly. On the contrary, our Q-VLM mines the cross-layer dependency of output distribution across layers, minimizing the block-wise discretization errors to avoid suboptimal quantization. We further optimize the visual encoder to disentangle the cross-layer dependency for fine-grained search space decomposition. As a result, our method outperforms ZeroQuant-V2 by 1.71 (79.79 vs. 78.08) in answering accuracy on ScienceQA dataset under 4-bit in LLaVA-7B model. Additionally, our method enhances inference speed, exceeding ZeroQuant-V2 by 1.2h (6.1h vs. 7.3h) due to utilizing stored rounding parameters instead of dynamic per-token quantization. The additional baseline provides a more comprehensive evaluation framework to highlight the strengths of our approach.
Model Quantization Method W8A8 W4A4
Accuracy Inference Time Accuracy Inference Time
LLaVA-7B ZeroQuant-V2 89.04 10.7h 78.08 7.3h
Q-VLM 89.58 8.3h 79.79 6.1h
LLaVA-13B ZeroQuant-V2 89.13 12.6h 78.81 9.7h
Q-VLM 89.81 11.2h 80.78 8.9h
Table 6: Comparisons with different quantization methods for 7B and 13B models across W6A6 and W4A4 bitwidth settings.
Appendix C Performance on other multi-modal architectures
We also explored the multi-modal architecture OpenFlamingo[3] to ensure the robustness and generalizability of our methods 7. We deploy our method on OpenFlamingo 3B model using Vizwiz and Hateful Memes[22] datasets, selecting bitwidths of 4 and 8 for quantized layers. Q-VLM designed in LLaVA-like architectures can be effectively adapted to cross-attention based VLMs due to the consistent core mechanism of cross-attention and the robust multimodal alignment capabilities pre-trained on large-scale vision-language pairs. Since OpenFlamingo is a cross-attention based VLM, exploiting cross-layer dependency is particularly suitable. Our method outperforms Q-LoRA by 1.22 (37.60 vs. 36.38) under 8-bit in OpenFlamingo-3B model. The advantage of our method becomes more obvious for 4-bit 3B LVLMs because quantization errors and cross-layer dependency play a more significant role in networks with low capacity. These results underscore the robustness and generalizability of our approach across different tasks, model architectures and datasets, demonstrating its effectiveness in diverse scenarios.
Dataset Shots FP 8bit 4bit
Q-LoRA Q-VLM Q-LoRA Q-VLM
Vizwiz 0 23.79 21.24 21.47 17.62 18.69
4 27.05 25.83 26.59 24.17 24.55
32 39.76 36.38 37.60 31.64 35.52
Hateful Memes 0 50.23 47.75 49.12 43.86 44.22
4 50.10 48.62 49.55 45.12 45.26
32 50.27 50.02 51.05 45.76 47.84
Table 7: Performance comparison on Vizwiz and Hateful Memes datasets across FP, 8bit, and 4bit quantization methods with different shot settings.